OpenStudy (nyasia):

A suit is on sale for 285$, which is 76% of the regular price. What is the regular price?

OpenStudy (mathstudent55):

The original price of the suit is 100% of the price. The discount was 76% of the original price. That means the amount paid is 100% - 76% = 34% of the original price.

OpenStudy (anonymous):

285 x 76%=216.60+285=501.60

OpenStudy (mathstudent55):

Just one interesting but crucial point. Do you mean "76% of the regular price" or "76% off the regular price" ?

OpenStudy (anonymous):

rather crucial F i would say...

OpenStudy (mathstudent55):

What I wrote above is based on interpreting the information as a discount of 76% off the regular price.

OpenStudy (nyasia):

Its 76 of the regular price

OpenStudy (mathstudent55):

Then disregard what I wrote above.

OpenStudy (mathstudent55):

Let's start again.

OpenStudy (mathstudent55):

The price on sale is $285 which is 76% of the regular price. Then the question is: $285 is 76% of what number?

OpenStudy (mathstudent55):

Since the regular price, x, (which we are looking for) is 100% of the price, we can set up a proportion: x it to 100% as 285 is to 76% \(\dfrac{x}{100} = \dfrac{285}{76} \) Cross multiply and solve for x.

OpenStudy (nyasia):

Thank You !

OpenStudy (mathstudent55):

You're welcome.